OpenStudy (anonymous):

A sum of money is invested at 6% simple interest. $2500 more than that amount is invested at 8% simple interest. If the total interest earned in one year from the two investments is $1068, how much was invested at 8%?

OpenStudy (anonymous):

(6x/100)+(8/100(x+2500))=1068 14x+20000=106800 14x=86800 x=6200 so the money invested at 8%=6200+2500=8700